# Enhancing Android Malware Detection: The Influence of ChatGPT on Decision-centric Task

### v1

@@ 1.4. Contributions
  • We are the first to comprehensively study the influence of ChatGPT on Android malware detection. Our research highlights that existing solutions lack interpretability, while ChatGPT excels in analysis and comprehension.
  • Our designed experiments and questionnaires reveal user and developer preferences for large language models like ChatGPT due to its high interpretability and powerful analysis capabilities.
  • Our research proposes a novel perspective for shaping the future of large language models for Android malware detection: alongside decision-making, more explanation.
The remainder of this article is structured as follows. Section 2 introduces the motivation and the relevant information on ChatGPT and Android malware detection. The proposed approach is presented in Section 3. Section 4 describes the experimental setup and evaluation metrics. Experimental results are shown in Section 5. The further analysis of the proposed approach is discussed in Section 6.2. Section 7 presents the related work. Finally, Section 8 provides a summary of the proposed approach in this paper.
@@ 2.1. Motivation
The development of Android malware detection has progressed significantly over the years, giving rise to numerous effective detection strategies. In theory, these strategies have demonstrated remarkable efficacy, but they still suffer dataset bias. For our study, we select six such strategies to conduct experiments. Therefore, we collected 67 instances of new malware introduced between July 1, 2023, and August 31, 2023. The results of their detection are presented in Table 1.
@@
From Table 1, it can be observed that the six schemes perform poorly in detecting unknown malicious applications. While these methods have demonstrated good results in training and testing datasets, a significant decline can be observed when detecting unfamiliar malware samples. This phenomenon raises our concerns regarding the reliability of the detection results. Meanwhile, these detection solutions have not provided detailed analysis and explanations. This means that we do not know the basis of their decision-making. Moreover, this raises our concerns about the reliability of these detection schemes. Consequently, we undertake a more in-depth exploration. As illustrated in Figure 1, we display the detection report of unknown application on the MalwareBazaar  [^2] https://bazaar.abuse.ch/  [^3] https://bazaar.abuse.ch/sample/44d0a81bdd9bbd3892f0ec49c430d60717be35e15a1a9ff3c2b4fddae58ad45c/ website. This site uploaded the application to eight popular antivirus tools. Among them, two tools deem the software harmless, but three label this unknown software as malicious, two consider it suspicious, and one does not provide any result. This indicates that even these advanced antivirus tools are unable to make a unanimous decision regarding an unknown malicious application. All of these tools provide their detection results, with some even offering detailed reports. However, these reports suffer from poor readability and require a wealth of specialized knowledge to comprehend. Moreover, these reports do not provide the rationale behind their decisions, nor do they offer explanations or analyses. Therefore, this demonstrates that both advanced detection schemes and popular antivirus tools have neglected the importance of explanation and analysis. This situation motivates us to conduct in-depth research to improve interpretability, thereby enhancing Android malware detection.
Meanwhile, we survey 101 practitioners to gather insights into their perspectives on current Android malware detection methods. We ask each participant 6 questions listed in Table 2. We design the survey in a semi-close-ended fashion where most questions have the opportunity to write down additional feedback. The participants’ roles include software engineers, developers, managers, etc. Table 3 shows the distribution of the participants over their profession and years of experience. We used the snowball sampling (Goodman, 1961) approach to select the participants. First, we picked 19 senior individuals (from our personal contacts) having at least 10 years of experience in the industry. We recorded their feedback and also requested them to invite appropriate candidates. The second batch of participants included 82 individuals ranging from developers to managers.
@@
All respondents indicate that they utilize antivirus software to safeguard their devices, suggesting that such software has become a fundamental necessity for individuals and businesses alike. Whether for personal or professional use, antivirus software enjoys considerable popularity, with individuals generally perceiving it as a crucial measure for ensuring system security. This demonstrates that respondents, regardless of their technical expertise, exhibit a notable degree of protection awareness.
@@
Next, we ask the participants about their reaction/experience of using antivirus software. The vast majority of respondents (91%) express satisfaction with the performance of their antivirus software, indicating that the majority of users believe it can effectively protect their systems. However, a notable minority (5%) report negative experiences, suggesting some level of dissatisfaction or a neutral attitude. Additionally, 4% of respondents indicate that their experience with this software is not positive and identify potential areas for improvement, particularly in functional performance and user experience.
@@
With regard to the comprehension of the detection reports, the distribution of respondents’ answers is relatively dispersed. Approximately 34.7% of respondents indicate that they could fully comprehend the report content, suggesting that these individuals possess a robust technical background or familiarity with the report’s subject matter. However, 11.9% of respondents indicate that their understanding of the report is limited, suggesting that the technical terminology or structure of the report may present a challenge. Of greater concern is that more than half of the respondents (53.4%) do not understand or are unclear about the report content, indicating that antivirus software reports may be too complex for ordinary users and lack simple and clear explanations.
@@
The survey results indicate that 95% of respondents had utilized ChatGPT, which is indicative of the pervasive adoption of this AI tool across technical and non-technical domains. This finding underscores the recognition of ChatGPT as a pivotal problem-solving instrument and its significant role in daily work activities. Conversely, only 2.9% of respondents report no usage of ChatGPT, which could be attributed to their professional requirements or a lack of familiarity with such tools.
@@
To ensure the interpretability of ChatGPT in SE real-world situations, we need to assess its ability to- i) interpret specific scenarios and contexts accurately; ii) generate appropriate information based on that. Hence, we provide ChatGPT with a specific question-answer set (about Glide) from Stack Overflow and ask it to explain how the library is used solely based on the provided information. In response to the question of whether ChatGPT could provide a more detailed explanation of the problem, 97% of respondents indicate that they find the explanation provided by ChatGPT to be sufficiently detailed and clear. This suggests that this cohort of users is content with their capacity to resolve issues. Nevertheless, 3% of respondents indicate a preference for more comprehensive explanations from ChatGPT, suggesting a need for more detailed information.
@@
In regard to the enhancement of malware detection, 69.3% of respondents indicate that the introduction of AI and machine learning technology would facilitate the improvement of detection capabilities. This finding suggests a high level of expectation for the integration of automation and intelligent technology in this domain. A 9.9% proportion of respondents express the desire to improve both the detection function and the software’s accuracy in identifying threats, indicating that existing functions could benefit from enhancement. Furthermore, 3% of respondents underscore the necessity for simplifying the user experience, proposing that a more intuitive interface and operational methodology would facilitate the software’s usability. The remaining 17.8% of respondents provide responses that are unclear, indicating that they may lack specific suggestions on how to improve malware detection or may be uncertain about their specific needs.
@@ 3.2. Feature Extraction
An APK file serves as the installation package for Android applications, encompassing both the program code and the essential resources for its execution. To decompile this APK file, we employ the reverse engineering tools, Apktool (Tumbleson and RyszardWiśniewski, 2023) and Androguard (Desnos, 2023). This process allows us to retrieve the AndroidManifest.xml file and the Class.dex file. The Class.dex file is then transformed into a legible Dalvik bytecode file, specifically the .smali format.
The AndroidManifest.xml serves as a configuration document, detailing aspects such as requisite permissions and the communicative components vital for the application’s operation. Analogously, each .smali file can be equated to a Java .class file in terms of its functionality.
According to previous research (Guerra-Manzanares et al., 2021), we extract 200 static features, encompassing elements like permissions, intent filters, and metadata. Further, upon installing the application on an actual device, we extract dynamic attributes, culminating in a list of 289 dynamic features, primarily characterized by system calls.
@@ 4. Experimental setup
To comprehensively assess the reliability of current models for data comprehension, we are currently conducting a series of experiments using the publicly accessible dataset, i.e., Kronodroid dataset (Guerra-Manzanares et al., 2021). Kronodroid dataset encompasses information collected from both simulator-generated data and actual devices. The real device data includes a total of 28,745 malicious samples, representing 209 distinct malware families, alongside 35,256 benign samples. Our data analysis encompasses the extraction of 289 dynamic features, specifically system calls, and 200 static features, including permissions, intent filters, and metadata. In our research, we randomly select 100 benign samples and 100 malicious samples from the Kronodroid dataset to form $dataset_{A}$ and conduct experiments and analyses. In addition, we also randomly select 3,000 benign samples and 3,000 malicious samples from the $Drebin$ (Arp et al., 2014) and AMD datasets (Wei et al., 2017) to form $dataset_{T}$, which is used for model training
The subsequent experiments are conducted on a personal computer integrated with 13th Gen Intel(R) Core(TM) i9-13900K 3.00 GHz and NVIDIA GeForce RTX 4090. The computer has 32GB of memory and 2TB of storage. The deep neural networks are implemented using Scikit-learn, and Pytorch.
@@ 5.1. Reliability of data understanding
In order to conduct a comprehensive comparison and analysis between ChatGPT and existing detection solutions, we invite ten experienced developers, each with more than 10 years of experience in Android, to participate in a questionnaire. The specific questions are as follows:
@@
Figure 9. Respondents’ feedback on reports of existing detection solutions. The vertical axis represents the scores, while the horizontal axis represents the interviewees, each numbered from 1 to 10.
Figure 10. Respondents’ feedback on reports of ChatGPT. The vertical axis represents the scores, while the horizontal axis represents the interviewees, each numbered from 1 to 10.
@@ 5.2. Influence and improvement
In contrast, compared to these decision-based detection solutions, ChatGPT, although unable to make explicit decisions, can provide a wealth of rich and detailed analysis and explanations. These analyses and explanations allow developers to gain deeper insights into the functionality used by the software and potential issues it may pose. Therefore, in order to improve and enhance detection solutions, we invite ten experienced developers, each with more than 10 years of experience in Android, to participate in a questionnaire. to participate in a carefully designed survey to further explore these aspects. We prepare four questions to gain a deeper understanding:
@@
Figure 11. Respondents’ feedback on potential enhancements of existing Android malware detection solutions. The vertical axis represents the scores, while the horizontal axis represents the interviewees, each numbered from 1 to 10.

### v2

@@ 1.4. Contributions
  • We conduct the first comprehensive study on the impact of ChatGPT in Android malware detection, identifying its strengths in interpretability and analysis compared to existing solutions.
  • Through experiments and user/developer surveys, we demonstrate the strong preference for large language models like ChatGPT due to their interpretability and analytical power.
  • We introduce a novel perspective for future Android malware detection: prioritizing explanation alongside decision-making to enhance usability and trust in large language models.
The remainder of this article is structured as follows. Section 2 introduces the motivation and the relevant information on ChatGPT and Android malware detection. The proposed approach is presented in Section 3. Section 4 describes the experimental setup and evaluation metrics. Experimental results are shown in Section 5. The further analysis of the proposed approach is discussed in Section 6. Section 7 presents the related work. Finally, Section 8 provides a summary of the proposed approach in this paper.
@@ 2.1. Motivation
The development of Android malware detection has progressed significantly over the years, giving rise to numerous effective detection strategies. In theory, these strategies have demonstrated remarkable efficacy. However, they still suffer from dataset bias, particularly selection bias and label bias, which stem from imbalanced and inconsistent datasets used during model training. To investigate this, we select six representative strategies for our study and conduct experiments to evaluate their performance. Therefore, we collected 67 instances of new malware introduced between July 1, 2023, and August 31, 2023, from the MalwareBazaar [^2] https://bazaar.abuse.ch/ website. This platform uploads newly discovered malware daily, providing direct download access to researchers and analysts. The results of their detection are presented in Table 1.
@@
From Table 1, it can be observed that the six schemes perform poorly in detecting unknown malicious applications. While these methods have demonstrated good results in training and testing datasets, a significant decline can be observed when detecting unfamiliar malware samples. This phenomenon raises our concerns regarding the reliability of the detection results. Meanwhile, these detection solutions have not provided detailed analysis and explanations. This means that we do not know the basis of their decision-making. Moreover, this raises our concerns about the reliability of these detection schemes. Consequently, we undertake a more in-depth exploration. As illustrated in Figure 1, we display the detection report of unknown application on the MalwareBazaar [^3] https://bazaar.abuse.ch/sample/44d0a81bdd9bbd3892f0ec49c430d60717be35e15a1a9ff3c2b4fddae58ad45c/ website. This site uploaded the application to eight popular antivirus tools. Among them, two tools deem the software harmless, but three label this unknown software as malicious, two consider it suspicious, and one does not provide any result. This indicates that even these advanced antivirus tools are unable to make a unanimous decision regarding an unknown malicious application. All of these tools provide their detection results, with some even offering detailed reports. However, these reports suffer from poor readability and require a wealth of specialized knowledge to comprehend. Moreover, these reports do not provide the rationale behind their decisions, nor do they offer explanations or analyses. Therefore, this demonstrates that both advanced detection schemes and popular antivirus tools have neglected the importance of explanation and analysis. This situation motivates us to conduct in-depth research to improve interpretability, thereby enhancing Android malware detection.
Meanwhile, we conducted a survey involving 101 practitioners to gather insights into their perspectives on current Android malware detection methods. Each participant was asked six questions, as listed in Table 2. The survey was designed in a semi-close-ended format, allowing participants to provide additional written feedback to supplement their answers.
The participants held diverse roles, including software engineers, developers, managers, and other professionals. Among them, 13 individuals were directly involved in the development of antivirus software, offering specialized insights into the field. The remaining participants worked in related areas such as mobile software development, operations, and maintenance. Table 3 provides a detailed breakdown of the participants’ professions and years of experience.
To select participants, we employed the snowball sampling approach (Goodman, 1961). Initially, we reached out to 19 senior professionals from our personal contacts, each with over 10 years of industry experience. After recording their feedback, we asked them to refer other suitable candidates for the study. This process led to the inclusion of an additional 82 participants, encompassing a broad spectrum of roles, from developers to managers. This approach ensured a diverse and experienced pool of respondents for the survey. The following is a detailed analysis of the questionnaire survey results.
@@
In response to the first question, ”Do you use antivirus software?”, all respondents reported utilizing antivirus software to protect their devices. This widespread adoption highlights the critical role of antivirus solutions as an essential tool for ensuring security, both for individuals and businesses. Whether for personal or professional use, antivirus software enjoys considerable popularity, with individuals generally perceiving it as a crucial measure for ensuring system security. This demonstrates that respondents, regardless of their technical expertise, exhibit a notable degree of protection awareness.
@@
Second question, we ask the participants about their reaction/experience of using antivirus software. The vast majority of respondents (91%) express satisfaction with the performance of their antivirus software, indicating that the majority of users believe it can effectively protect their systems. However, a notable minority (5%) report negative experiences, suggesting some level of dissatisfaction or a neutral attitude. Additionally, 4% of respondents indicate that their experience with this software is not positive and identify potential areas for improvement, particularly in functional performance and user experience.
@@
The third question posed to the interviewees is: ”How well do you understand the detection reports generated by antivirus software?” With regard to the comprehension of the detection reports, the distribution of respondents’ answers is relatively dispersed. Approximately 34.7% of respondents indicate that they could fully comprehend the report content, suggesting that these individuals possess a robust technical background or familiarity with the report’s subject matter. However, 11.9% of respondents indicate that their understanding of the report is limited, suggesting that the technical terminology or structure of the report may present a challenge. Of greater concern is that more than half of the respondents (53.4%) do not understand or are unclear about the report content, indicating that antivirus software reports may be too complex for ordinary users and lack simple and clear explanations.
@@
The survey results of the fourth question indicate that 95% of respondents had utilized ChatGPT, which is indicative of the pervasive adoption of this AI tool across technical and non-technical domains. This finding underscores the recognition of ChatGPT as a pivotal problem-solving instrument and its significant role in daily work activities. Conversely, only 2.9% of respondents report no usage of ChatGPT, which could be attributed to their professional requirements or a lack of familiarity with such tools.
@@
The fifth question is about the performance of ChatGPT. To ensure the interpretability of ChatGPT in SE real-world situations, we need to assess its ability to- i) interpret specific scenarios and contexts accurately; ii) generate appropriate information based on that. Hence, we provide ChatGPT with a specific question-answer set (about Glide) from Stack Overflow and ask it to explain how the library is used solely based on the provided information. In response to the question of whether ChatGPT could provide a more detailed explanation of the problem, 97% of respondents indicate that they find the explanation provided by ChatGPT to be sufficiently detailed and clear. This suggests that this cohort of users is content with their capacity to resolve issues. Nevertheless, 3% of respondents indicate a preference for more comprehensive explanations from ChatGPT, suggesting a need for more detailed information.
@@
The last question posed to the interviewees is: ”What can be the ways to enhance existing forms of malware detection?” In regard to the enhancement of malware detection, 69.3% of respondents indicate that the introduction of AI and machine learning technology would facilitate the improvement of detection capabilities. This finding suggests a high level of expectation for the integration of automation and intelligent technology in this domain. A 9.9% proportion of respondents express the desire to improve both the detection function and the software’s accuracy in identifying threats, indicating that existing functions could benefit from enhancement. Furthermore, 3% of respondents underscore the necessity for simplifying the user experience, proposing that a more intuitive interface and operational methodology would facilitate the software’s usability. The remaining 17.8% of respondents provide responses that are unclear, indicating that they may lack specific suggestions on how to improve malware detection or may be uncertain about their specific needs.
@@ 3.2. Feature Extraction
The APK file, which functions as the installation package for Android applications, is analyzed using reverse engineering tools such as Apktool (Tumbleson and RyszardWiśniewski, 2023) and Androguard (Desnos, 2023). Apktool decompiles the APK into its core resources and bytecode, extracting critical components like the AndroidManifest.xml file and converting the Class.dex file (containing Dalvik Executable bytecode) into a human-readable .smali format. This enables detailed examination of API calls, strings, and control flows. In contrast, Androguard offers a robust framework for static analysis, supporting disassembly, decompilation, and in-depth bytecode inspection. It complements Apktool by facilitating advanced operations such as cryptographic signature analysis and comprehensive behavioral insights.
The extracted AndroidManifest.xml file is parsed to collect static features such as declared permissions, intent filters, and metadata, providing valuable information about the app’s functionalities and interactions with the Android system. Additionally, the Class.dex file, once converted to .smali format, supports an in-depth exploration of the app’s logic, helping uncover potential malicious behaviors embedded within its control flows or API usage patterns.
For dynamic feature extraction, applications are executed in a controlled emulator environment where runtime behaviors are monitored. This includes tracking system calls, interactions with the Android system, and other runtime characteristics.
When obfuscated APK files are encountered, unpacking tools (Xue et al., 2017; Zhang et al., 2015) are employed to recover the underlying files. Features extracted from these samples are cross-verified against dataset-provided features to ensure their accuracy.
@@ 4. Experimental setup
To comprehensively assess the reliability of current models for data comprehension, we conducted a series of experiments using publicly accessible datasets. Specifically, we utilized the Kronodroid dataset (Guerra-Manzanares et al., 2021), which includes data collected from both emulator environments and actual devices. The dataset comprises 28,745 malicious samples across 209 distinct malware families and 35,256 benign samples. For our analysis, we focus on extracting 289 dynamic features (e.g., system calls) and 200 static features (e.g., permissions, intent filters, and metadata). In our research, we randomly selected 100 benign samples and 100 malicious samples from the Kronodroid dataset, enabling rapid experimentation and analysis. Additionally, we constructed $dataset_{T}$ by randomly selecting 3,000 benign samples and 3,000 malicious samples from the Drebin (Arp et al., 2014) and AMD datasets (Wei et al., 2017). Unlike prior studies that emphasize malware family classification, our research exclusively focuses on the binary classification of benign versus malicious applications, disregarding family information throughout dataset construction, model training, and detection processes. For obfuscated samples, we employed unpacking systems (Xue et al., 2017; Zhang et al., 2015) to recover original files, verified the extracted features against dataset-provided features, and ensured the accuracy of the analysis.
We construct two distinct datasets to support our analysis. The $dataset_{A}$, containing 200 samples (100 benign and 100 malicious), is primarily used to analyze ChatGPT’s performance and conduct comparative experiments. We adopt this size based on the dataset configurations commonly used in related research (Ali-Gombe et al., 2018; Abraham et al., 2015). Its main purpose is to explore ChatGPT’s ability to analyze malware rather than optimizing detection performance, and its size also helps minimize computational overhead, fitting within GPT-4’s usage constraints. The $dataset_{D}$, with 6,000 samples (3,000 benign and 3,000 malicious), is used for training traditional detection models, reflecting typical configurations in related research (Li et al., 2022b). The $dataset_{D}$ allows for better model training and generalization by covering a wider range of software behaviors. By using both datasets, we leverage the strengths of each: the smaller one enables quick analysis, while the larger one supports more comprehensive training and validation. The $dataset_{A}$ is utilized for the analysis and testing of ChatGPT, while $dataset_{D}$ is employed for the training and testing of traditional models. For a comprehensive overview of the experiments and results, please refer to Section 5.1.
The subsequent experiments are conducted on a personal computer integrated with 13th Gen Intel(R) Core(TM) i9-13900K 3.00 GHz and NVIDIA GeForce RTX 4090. The computer has 32GB of memory and 2TB of storage. The deep neural networks are implemented using Scikit-learn, and Pytorch. Meanwhile, the ChatGPT 4 model is employed for analysis in subsequent experiments.
@@ 5.1. Reliability of data understanding
In this section, we begin by verifying the effectiveness of existing detection solutions and identifying their limitations (Section 5.1.1). We conduct this experiment using two datasets: $dataset_{T}$ and $dataset_{A}$. The $dataset_{T}$ is divided into a training set and a test set, with the training set used to train the models and the test set used to evaluate their performance. Subsequently, we apply the trained models to $dataset_{A}$ to assess their performance on previously unseen data, providing insights into the generalizability of the detection schemes. Next, we assess the analysis and explanation capabilities of ChatGPT (Section 5.1.2). Finally, we compare the performance of traditional models with that of ChatGPT (Section 5.1.3).
5.1.1. Performance of Existing Methods
@@ 5.1.2. Performance of ChatGPT
5.1.3. Human Evaluation
In order to conduct a comprehensive comparison and analysis between ChatGPT and existing detection solutions, we invite ten experienced developers, each with more than 10 years of experience in Android, to participate in a questionnaire. All ten respondents are engaged in Android mobile software development, which aligns closely with the focus of our study. Their work involves various aspects of the Android ecosystem, including coding, testing, and maintaining mobile applications. This hands-on experience ensures they possess in-depth knowledge of Android development frameworks, tools, and best practices. The specific questions are as follows:
@@ 5.1.3. Human Evaluation
Figure 9. Respondents’ feedback on reports of existing detection solutions. The vertical axis displays the scores, ranging from 0 to 5, while the horizontal axis represents the respondents, numbered sequentially from the first to the tenth.
Figure 10. Respondents’ feedback on reports of ChatGPT. The vertical axis represents the score, while the horizontal axis represents the respondents. The vertical axis displays the scores, ranging from 0 to 5, while the horizontal axis represents the respondents, numbered sequentially from the first to the tenth.
@@
The survey responses offer insightful perspectives on Android malware detection solutions and the preferences of experienced developers. Most respondents, who frequently use Android-based devices both professionally and personally, demonstrate a high reliance on antivirus software. This underscores the importance of effective malware detection tools in real-world applications.
In contrast, ChatGPT’s reports were universally praised for their clarity, detail, and comprehensive analysis, making them more intuitive and user-friendly. However, respondents also noted its limitation in providing final decisions or judgments, which is a significant drawback for practical application in malware detection. Despite this, the developers overwhelmingly preferred ChatGPT for its readability, comfort, and overall user-friendliness, reinforcing the importance of clear and understandable explanations in building user trust. Overall, while existing Android malware detection models are widely used, their lack of transparency and interpretability remains a key challenge. ChatGPT’s strength lies in its ability to offer detailed explanations, making it a valuable tool for developers, though further development is needed to enable it to provide definitive decisions.
@@ 5.2. Influence and improvement
In contrast, compared to these decision-based detection solutions, ChatGPT, although unable to make explicit decisions, can provide a wealth of rich and detailed analysis and explanations. These analyses and explanations allow developers to gain deeper insights into the functionality used by the software and potential issues it may pose. Therefore, in order to improve and enhance detection solutions, we invite ten experienced developers to participate in a questionnaire. Six of the ten respondents are actively involved in antivirus software development, contributing their specialized expertise in identifying and mitigating security threats. Three respondents are engaged in artificial intelligence development, including predictive analytics, machine learning, and intelligent automation. The remaining respondent focuses on Android software development. Their expertise adds valuable context to the analysis, particularly in examining the practical implications of malware detection systems within the Android environment. We prepare four questions to gain a deeper understanding:
@@
Figure 11. Respondents’ feedback on potential enhancements of existing Android malware detection solutions. The vertical axis displays the scores, ranging from 0 to 5, while the horizontal axis represents the respondents, numbered sequentially from the first to the tenth.
@@
The survey results reveal significant insights into the current state of Android malware detection solutions and highlight key areas for improvement. While all respondents have experience with detection tools, the majority find them challenging to use, with only 20% rating them as user-friendly. The remaining 80% encounter difficulties such as complex interfaces, a lack of user-friendly documentation, and time-consuming setup processes, which create barriers to accessibility, especially for non-experts.
Regarding ChatGPT, all respondents acknowledged its strengths in analytical capabilities and ease of use, particularly its ability to provide clear, detailed explanations in a conversational format. However, the inability of ChatGPT to make definitive decisions was identified as a major drawback, as decision-making is crucial in malware detection tasks.
In conclusion, the survey highlights the need for more user-friendly and interpretable malware detection tools that not only simplify the user experience but also provide clear insights into decision-making processes. ChatGPT shows promise due to its analytical strengths and ease of use, but its inability to make concrete decisions indicates the need for further development to make it a more practical tool for malware detection. The future of Android malware detection lies in integrating robust explanatory features with actionable decision-making capabilities, ensuring both accessibility and reliability for developers.